OpenStudy (anonymous):

A line has a slope of -5 and passes through the point (2, 2). An equation for this line is

OpenStudy (anonymous):

You want to find the equation for a line that passes through the point (2,2) and has a slope of -5. First of all, remember what the equation of a line is: y = mx+b Where: m is the slope, and b is the y-intercept To start, you know what m is; it's just the slope, which you said was -5. So you can right away fill in the equation for a line somewhat to read: y=-5x+b. Now, what about b, the y-intercept? To find b, think about what your (x,y) point means: (2,2). When x of the line is 2, y of the line must be 2. Because you said the line passes through this point, right? Now, look at our line's equation so far: . b is what we want, the -5 is already set and x and y are just two "free variables" sitting there. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the the point (2,2). So, why not plug in for x the number 2 and for y the number 2? This will allow us to solve for b for the particular line that passes through the point you gave!. (2,2). y=mx+b or 2=-5 × 2+b, or solving for b: b=2-(-5)(2). b=12. The equation of the line that passes through the point (2,2) with a slope of -5 is y=-5x+12

OpenStudy (anonymous):

Point-Slope form: y - y1 = m(x - x1) y - 2 = -5(x - 2) y - 2 = -5x + 10 +2 y = -5x + 12 is your answer:)

OpenStudy (anonymous):

that cat be righty=x2−2 y=−x2−1 y=2x+1 y=−2x−1